OpenStudy (anonymous):

a/b =10a what number should B be?

OpenStudy (anonymous):

a/b =10a /1 1/b = 10/1

OpenStudy (anonymous):

why B became 1?

OpenStudy (anonymous):

a/b =10a Try to get similar terms on one side (a one side, b one side) Right? If you multiply both sides by b first you get a = 10a*b (* is multiplification) You got rid of fractions first because then you'd know you could see what to divide by to get only similar terms to one side (a one side, b the other side) So now you see that you can divide by a to get: 1 on left side (a/a) 10b on one side ((10a*b) /a = 10*b = 10b) so 1 on left, 10b on right 1 = 10b Now you just divide both sides by 10 to find what is the value of 1b, not whwat is the value of 10b so 1/10 = b